OpenStudy (anonymous):

use the shell method to find the volume bounded by y= 2sqrt(x), and y=6 rotated around the y-axis

OpenStudy (anonymous):

Do you know how to get started with this problem? You can start by drawing the graph.

OpenStudy (anonymous):

yes i have the graph already i just don't know how to get the answer

OpenStudy (anonymous):

Ok, so how far into the problem do you understand it?

OpenStudy (anonymous):

^or what part are you having problems with

OpenStudy (anonymous):

I think I started off like this \[2\pi \int\limits_{0}^{6}x(2\sqrt{x})dx\]

OpenStudy (anonymous):

but I am not sure

OpenStudy (anonymous):

and if its correct I dont know how to continue

OpenStudy (anonymous):

Have you drawn the two functions? That's usually a good start.

OpenStudy (anonymous):



OpenStudy (anonymous):

Your integral is almost correct, make sure to check at what point both functions intersect.

OpenStudy (anonymous):